### v1

# Schauder Bases for $C[0,1]$ Using ReLU, Softplus and Two Sigmoidal Functions

Abstract
We construct four Schauder bases for the space $C[0,1]$, one using ReLU functions, another using Softplus functions, and two more using sigmoidal versions of the ReLU and Softplus functions. This establishes the existence of a basis using these functions for the first time, and improves on the universal approximation property associated with them.
1 Introduction
Functions in spaces such as $C[0,1]$ with the supremum norm, $L^{2}[0,1]$, and $L^{2}(\mathbb{R})$ can be approximated, or in Kolmogorov’s sense represented, using finite linear combinations of the form:
$\sum_{i=1}^{N}\alpha_{i}\sigma(w_{i}x+b_{i})$ (1)
Expressions like (1) resemble single hidden-layer neural networks. The component functions $\sigma(w_{i}x+b_{i})$, also known as plane waves, ridge functions, or sigmoids, have been widely applied in fields such as finance, data analysis, statistics, and medical imaging [Ismailov2021].
Although originally linked to Kolmogorov’s representation theorem [Kolmogorov1957, Sprecher1965], this formulation has found broader utility in function approximation, as neatly delineated by Sprecher [Sprecher2021]. This shift is seen in Cybenko’s universal approximation theorem [Cybenko1989], where $\sigma$ is assumed to be a fixed sigmoidal function. In the Kolmogorov-Sprecher representational approach, the function $\sigma$ depends on the specific target function $f$. In contrast, Cybenko’s approximation perspective assumes $\sigma$ is fixed and independent of $f$.
Our work extends beyond these two perspectives by constructing infinite series representations using fixed activation functions. Specifically, we present two Schauder bases for $C[0,1]$: one based on the widely used ReLU function (Theorem 1) and the other on its smooth variant, the Softplus function (Theorem 3).
In Banach space literature, functions similar to ReLU have appeared in basis constructions, such as the Schauder hat and restricted hat functions [Semadeni1982, p.28]. However, the restricted hat, which is essentially a ReLU, is only used to describe the boundary behavior of a Schauder hat function, and not to construct a full basis. The ReLU and Softplus functions have not been previously employed for full basis constructions, possibly because their development was motivated by later applications in neural networks.
Within ridge function literature there is work on a universal sigmoidal function independent of the target function $f$. The existence of such a universal function for target functions in $C(\mathbb{R})$, with a prescribed approximation error, and a prescribed number of neurons is guaranteed by a theorem of Maiorov and Pinkus [Ismailov2021, p.158], and an algorithmic construction is provided by Ismailov[Ismailov2021, p.164]. Our basis construction (Theorem 1) seemingly provides such a universal function for $C[0,1]$, but this is misleading as the number of neurons will depend on the level of approximation desired. But again, the ability to approximate arbitrarily well with a fixed number of neurons depends on the use of wild and pathological functions, as Sprecher notes in [Sprecher2021], or the use of intricate algorithms as in Ismailov’s smooth, almost monotone construction. In contrast, we use simple and standard functions like the ReLU in our basis constructions.
In the context of neural networks, having a basis offers certain potential benefits compared to density results like those in [Cybenko1989]. For example, a basis allows for structured initialization in neural network training. Consider an infinite expansion of the form:
$f(x)=\sum_{i=1}^{\infty}\alpha_{i}\sigma_{i}(x),\quad\text{where}\sigma_{i}(x)=\sigma(w_{i}x+b_{i})$
This expansion is to be interpreted in the sense of a Schauder basis, as discussed in the next section. A finite truncation yields:
$f(x)\approx f_{M}(x)=\sum_{i=1}^{M}\alpha_{i}\sigma_{i}(x)$
Here, $f_{M}(x)$ can be seen as a neural network of width $M$. To improve the approximation, one could consider a wider network $f_{N}(x)$ with $N>M$, retaining both the functions $\sigma_{i}$ and coefficients $\alpha_{i}$ for $i\leq M$, and only learning the new coefficients for $i>M$. This reuse of parameters from a narrower model is theoretically justified only when an explicit basis is available. In contrast, density results offer no such structure and require retraining from scratch.
Another possible advantage of using a basis is that it ensures unique expansions, and thus a unique global minimmum during training. Additionally, convergence under the $C[0,1]$ supnorm topology may impact outlier behavior during model inference, though it remains unclear what the impact will be. It is well known that small perturbations of the input can have a drastic effect on neural network output, and it is possible that using supnorm topology provides some control on this behavior compared to the use of $L_{2}$ loss functions.
From a theoretical standpoint, Cybenko [Cybenko1989] notes that completeness results typically fall into two broad categories: those related to Weierstrass’s theorem on polynomial density and those based on Wiener’s translation-invariant systems. In a way, our results incorporate aspects of both, much like Schauder’s original basis from 1927.
Our construction, like Weierstrass’s result, is situated within $C[0,1]$ and relies on a discrete bump function - the Schauder hat. At the same time, it employs scaled and shifted versions of a single activation function, akin to the translation-invariant methods of Wiener. This use of scaling and shifting is common in $L_{2}[0,1]$ literature, and in wavelet theory by extension.
2 Preliminaries
A countable sequence $\{x_{n}\}$ in a Banach space $X$ is a basis for $X$ if for all $x$ in $X$ there exist unique scalars $a_{n}(x)$ such that
$x=\sum_{n=1}^{\infty}a_{n}(x)x_{n}$ (2)
where the above series converges in the norm of $X$ [Heil2010]. A normalized basis is a basis $\{x_{n}\}$ with $||x_{n}||=1$ for all $n$. A Schauder basis for $X$ is a basis for $X$ where the $a_{n}$’s are continuous linear functionals. There are other definitions of a Schauder basis that involve uncountable index sets, but we restrict ourselves to the above countable version. As it turns out, any basis for a Banach space is a Schauder basis. When $X=C[0,1]$ we use the standard supnorm topology. Since $C[0,1]$ is not a Hilbert space there is no notion of an inner product, and the linear functionals $a_{n}(x)$ are the closest approximation to the usual coordinates $\langle x_{n},x\rangle$ based on an inner product .,$\langle$.,$\rangle$ . , . ⟩.
The ReLU function $r(x)$ is defined as follows:
$r(x)=\left\{\begin{array}[]{ll}0&\mbox{if }x<0,\\ x&\mbox{if }x\geq 0.\end{array}\right.$ (5)
The parameterized Softplus function defined as follows [Dugan2000]:
$p_{a}(x)=\frac{\ln(1+e^{ax})}{a}.$ (6)
A function $\sigma:\mathbb{R}\to\mathbb{R}$ is called sigmoidal if
$\lim_{x\to\infty}\sigma(x)=1$ (7)
$\lim_{x\to-\infty}\sigma(x)=0$ (8)
Schauder’s original basis functions for $C[0,1]$ ([Heil2010, p.142]) are defined by the ordered set $\mathcal{S}=\{s_{n,k}|n\in\mathbb{N}\cup\{0\},k\in\{0,1,\ldots,2^{n}-1\}\}$ under dictionary ordering, where
$s_{n,k}(x)=\left\{\begin{array}[]{ll}1&\mbox{if }x=\frac{k+\frac{1}{2}}{2^{n}},\\ \text{linear}&\mbox{on }[\frac{k}{2^{n}},\frac{k+\frac{1}{2}}{2^{n}}]\text{ and }[\frac{k+\frac{1}{2}}{2^{n}},\frac{k+1}{2^{n}}],\\ 0&\text{otherwise.}\end{array}\right.$ (12)
Given $f\in C[0,1]$ let us denote its basis expansion by:
$f=\alpha_{0}\chi_{[0,1]}+\alpha_{1}s_{1}+\sum_{k<2^{n}}\alpha_{n,k}s_{n,k}.$ (13)
Besides $\alpha_{n,k}s_{n,k}$, the Schauder basis expansion contains two additional terms $\alpha_{0}\chi_{[0,1]}$ and $\alpha_{1}s_{1}$. These involve two basis functions, namely the characteristic function $\chi_{[0,1]}$ and the linear function $s_{1}(x)=x$. We assume standard ordering for the Schauder basis, which is essentially a dictionary ordering of $(n,k)$. The ordering is important for convergence since the basis is conditional. We will now look at a basis construction for $C[0,1]$ based on the above ReLU function $r(x)$.
3 ReLU Basis
We now construct a Schauder basis for $C[0,1]$ using the ReLU function $r(x)$ and two auxiliary functions. In particular, we prove the following theorem:
Theorem 1.
The basis functions $\chi_{[0,1]}(x)$, $s_{1}(x)=x$, $r(2^{n}x-k)$ and $r(2^{n}x-(k+\frac{1}{2}))$ form a Schauder basis for $C[0,1]$. In particular, borrowing $\alpha_{n,k}$ from the Schauder basis expansion (12), and setting $\alpha_{n,-1}=0$, we have
$f=\alpha_{0}\chi_{[0,1]}+\alpha_{1}s_{1}+\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\{(\alpha_{n,k}+\alpha_{n,k-1})r(2^{n}x-k)-2\alpha_{n,k}r(2^{n}x-(k+\frac{1}{2}))\}.$ (14)
where the coefficient functionals $\alpha_{0}$, $\alpha_{1}$, $(\alpha_{n,k}+\alpha_{n,k-1})$ and $-2\alpha_{n,k}$ are all bounded.
Please note that we have simplified notation by implicitly restricting our basis functions to $[0,1]$, and we will continue to do so. i.e. we will write $f(x)$ for $f(x)|_{[0,1]}$.
We begin our proof with the following lemma:
Lemma 1.
Define $t_{n,k}$ as follows:
$t_{n,k}(x)=r(2^{n}x-k)-2r(2^{n}x-(k+\frac{1}{2}))+r(2^{n}x-(k+1)).$ (15)
. We claim that $t_{n,k}(x)=s_{n,k}(x)$ for all $x\in\mathbb{R}$.
Proof.
Let $x=\frac{k+\delta}{2^{n}}$. Then,
$t_{n,k}$ $=r(2^{n}\frac{k+\delta}{2^{n}}-k)-2r(2^{n}\frac{k+\delta}{2^{n}}-(k+\frac{1}{2}))+r(2^{n}\frac{k+\delta}{2^{n}}-(k+1))$
$=r(\delta)-2r(\delta-\frac{1}{2})+r(\delta-1).$
We will now establish equality of $t_{n,k}$ and $s_{n,k}$ in all the four cases listed in (12).
For $x=\frac{k+\frac{1}{2}}{2^{n}}$, we have $\delta=\frac{1}{2}$
$t_{n,k}$ $=r(\delta)-2r(\delta-\frac{1}{2})+r(\delta-1)$
$=r(\frac{1}{2})-2r(0)+r(\frac{-1}{2})$
$=r(\frac{1}{2})$
$=1.$
For $x\in[\frac{k}{2^{n}},\frac{k+\frac{1}{2}}{2^{n}}]$, $\delta\in[0,\frac{1}{2}]$. Thus,
$t_{n,k}=r(\delta)=\delta=2^{n}x-k,$
which is linear in $x$.
Similarly, for $x\in[\frac{k+\frac{1}{2}}{2^{n}},\frac{k+1}{2^{n}}]$, we have $\delta\in[\frac{1}{2},1]$. So,
$t_{n,k}$ $=r(\delta)-2r(\delta-\frac{1}{2})+r(\delta-1)$
$=\delta-2(\delta-\frac{1}{2})+0$
$=1-\delta=\frac{1}{2}-(2^{n}x-k)$
$=k+1-2^{n}x,$
which is also linear in $x$.
Next for $x<\frac{k}{2^{n}}$, $\delta\leq 0$. Thus,
$t_{n,k}$ $=r(\delta)-2r(\delta-\frac{1}{2})+r(\delta-1)$
$=0$
Finally, for $x>\frac{k+1}{2^{n}}$, $\delta>1$ and
$t_{n,k}$ $=r(\delta)-2r(\delta-\frac{1}{2})+r(\delta-1)$
$=\delta-2(\delta-\frac{1}{2})+(\delta-1)$
$=0.$
We have $t_{n,k}(x)=s_{n,k}(x)$ in all the cases considered, and thus $t_{n,k}(x)=s_{n,k}(x)$ for all $x\in\mathbb{R}$.
We will now prove Theorem 1.
Proof.
$f$ $=\alpha_{0}\chi_{[0,1]}+\alpha_{1}s_{1}+\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\alpha_{n,k}s_{n,k}$
$=\alpha_{0}\chi_{[0,1]}+\alpha_{1}s_{1}+\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\alpha_{n,k}t_{n,k}$
$=\alpha_{0}\chi_{[0,1]}+\alpha_{1}s_{1}+\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\alpha_{n,k}\{r(2^{n}x-k)-2r(2^{n}x-(k+\frac{1}{2}))+r(2^{n}x-(k+1))\}$
$=\alpha_{0}\chi_{[0,1]}+\alpha_{1}s_{1}+\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\{\alpha_{n,k}r(2^{n}x-k)-2\alpha_{n,k}r(2^{n}x-(k+\frac{1}{2}))+\alpha_{n,k}r(2^{n}x-(k+1))\}$
$=\alpha_{0}\chi_{[0,1]}+\alpha_{1}s_{1}+\sum_{n=0}^{\infty}\{\alpha_{n,0}r(2^{n}x)+\sum_{k=1}^{2^{n}-1}\{(\alpha_{n,k}+\alpha_{n,k-1})r(2^{n}x-k)-2\alpha_{n,k}r(2^{n}x-(k+\frac{1}{2}))\}\}$ ($\ast$)
$=\alpha_{0}\chi_{[0,1]}+\alpha_{1}s_{1}+\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\{(\alpha_{n,k}+\alpha_{n,k-1})r(2^{n}x-k)-2\alpha_{n,k}r(2^{n}x-(k+\frac{1}{2}))\}$
where, for convenience, we have set $\alpha_{n,-1}=0$ in the last step. Notice that ($\ast$3) does not involve any rearrangement of terms. The grouping of $(\alpha_{n,k}+\alpha_{n,k-1})$ involves only associativity, and no commutativity. In particular, the conditional convergence of the earlier series and that of ($\ast$3) are equivalent.
Finally, given that the coefficient functionals $\alpha_{n,k}$ are bounded, the coefficients of the ReLU expansion, namely, $\alpha_{0}$, $\alpha_{1}$, $\alpha_{n,k}+\alpha_{n,k-1}$ and $-2\alpha_{n,k}$ are all bounded as well. This establishes that the sequence of functions $\chi_{[0,1]}(x)$, $s_{1}(x)=x$, $r(2^{n}x-k)$ and $r(2^{n}x-(k+\frac{1}{2})$ form a Schauder basis. ∎
Remarks: Let $t(x)$ be defined as follows to be the discrete second derivative of $r(x)$:
$t(x)=r(x)-2r(x-\frac{1}{2})+r(x-1).$ (16)
$t(x)$ is a triangular bump or hat function which is the essential building block of the original basis of Schauder. Letting $d(x)$ denote the expression $r(x)-r(x-\frac{1}{2})$, we can see that $t(x)$ matches the expression $d(x)-d(x-\frac{1}{2})$. Here $d(x)$ can be interpreted as the first discrete derivative of $r(x)$, and $t(x)$ as the corresponding first discrete derivative of $d(x)$, or in effect the second discrete derivative of $r(x)$.
The subscripted functions $r_{n,k}(x)$, $d_{n,k}(x)$ and $t_{n,k}(x)$ are dyadically scaled and shifted versions of $r(x)$, $d(x)$ and $t(x)$, for instance $t_{n,k}(x)=t(2^{n}x-k)$. Correspondingly, $d_{n,k}(x)$ is the first discrete derivative of $r_{n,k}(x)$ and $t_{n,k}(x)$ is the first discrete derivative of $d_{n,k}(x)$, or in effect the second discrete derivate of $r_{n,k}(x)$. It is well known that the Haar basis elements represent the first derivative of the Schauder basis elements. The current construction shows that the Schauder basis elements represent the first discrete derivative of $d_{n,k}$ and the second discrete derivative of the ReLU basis elements $r_{n,k}$.
$d_{n,k}(x)$ is a continuous sigmoidal function, and thus dense in $C[0,1]$ as per Cybenko [Cybenko1989]. Going beyond Cybenko’s result, $d_{n,k}$ can be used to assemble a basis as shown in the following theorem:
Theorem 2.
The functions $d_{n,k}(x)$ with $n\geq 0$ and $k\leq 2^{n}-1$ along with the auxiliary functions $\chi_{[0,1]}(x)$ and $s_{1}(x)=2$ forms a Schauder basis for $C[0,1]$ with expansions of the following form for all $f$ in $C[0,1]$:
$f=\alpha_{0}\chi_{[0,1]}+\alpha_{1}s_{1}+\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\alpha_{n,k}d_{n,k}-\alpha_{n,k}d_{n,k-\frac{1}{2}}$ (17)
where the coefficient functionals are all bounded.
Proof.
We start with the basis expansion of $f$ using the basis elements $t_{n,k}$ where $\alpha_{j}$ and $\alpha_{n,k}$ are borrowed from the Schauder basis expansion (12).
$f$ $=\alpha_{0}\chi_{[0,1]}+\alpha_{1}s_{1}+\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\alpha_{n,k}t_{n,k}$
$=\alpha_{0}\chi_{[0,1]}+\alpha_{1}s_{1}+\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\alpha_{n,k}(d_{n,k}-d_{n,k-\frac{1}{2}})$
$=\alpha_{0}\chi_{[0,1]}+\alpha_{1}s_{1}+\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\alpha_{n,k}d_{n,k}-\alpha_{n,k}d_{n,k-\frac{1}{2}}$
This concludes the proof since the coefficient functionals $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{n,k}$ are all known to be bounded. ∎
We note a general principle at play here with regards to first and second discrete derivatives. In particular, if $f_{j}\in X$ form a Schauder basis for Banach space $X$, and if $f_{j}$ are the first discrete derivatives of $g_{j}$ and the second discrete derivatives of $h_{j}$, then $g_{j}$ and $h_{j}$ form Schauder bases for $X$ as well. One may argue that the basis property of $h_{j}$ follows from the basis property of $g_{j}$ by induction, but an examination of the above proofs shows that some care is required to avoid double counting as seen in the functional $(\alpha_{n,k}+\alpha_{n,k-1})$.
As noted earlier, $d_{n,k}$ forms a basis, but $d(x)$ is not a universal sigmoidal function for $C[0,1]$. Unlike the construction of Ismailov [Ismailov2021, p.164] for $C(\mathbb{R})$, the number of terms in a series truncation increases with the desired level of approximation.
4 Softplus Basis
Schauder bases posses some stability properties wherein the basis property holds even if each element is perturbed slightly. We use this stability property to perturb the ReLU basis and obtain a basis using Softplus functions. In particular, we will prove the following theorem:
Theorem 3.
The basis functions $\chi_{[0,1]}(x)$, $s_{1}(x)=x$, $p_{a(n,k)}(2^{n}x-k)$ and $p_{a(n,k)}(2^{n}x-(k+\frac{1}{2}))$ form a Schauder basis for $C[0,1]$. In particular, given $f$ in $C[0,1]$, we have the following basis expansion with $a(n,k)=4\ln 2\cdot 2K\cdot 2^{2n+2}$
$f=\gamma_{0}\chi_{[0,1]}+\gamma_{1}s_{1}+\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\{\gamma_{n,k}p_{a(n,k)}(2^{n}x-k)+\psi_{n,k}p_{a(n,k)}(2^{n}x-(k+\frac{1}{2})\}.$ (18)
where the coefficient functionals $\gamma_{0}$, $\gamma_{1}$, $\gamma_{n,k}$ and $\psi_{n,k}$ are all bounded.
We start by recalling some stability properties of Schauder bases. First, we have the notion of a basis constant whose existence is asserted in the following classical theorem [Linden1977].
Theorem 4.
Let $\{x_{n}\}$ be a Schauder basis of a Banach Space $X$. Then the projections $P_{n}:X\to X$ defined by $P_{n}(\sum_{i=1}^{\infty}a_{i}x_{i})=\sum_{i=1}^{n}a_{i}x_{i}$ are bounded linear operators and $\sup_{n}||P_{n}||<\infty$. The supremum $K=\sup_{n}||P_{n}||$ is called the basis constant of $\{x_{n}\}$.
We then have the following stability property [Linden1977, Prop. 1.a.9]:
Theorem 5.
Let $\{x_{n}\}$ be a normalized Schauder basis of a Banach Space $X$ with basis constant $K$. If $\{y_{n}\}$ is a sequence of vectors in $X$ such that $\sum_{i=1}^{\infty}||x_{n}-y_{n}||<\frac{1}{2K}$, then $\{y_{n}\}$ is also a Schauder basis of $X$.
To prove that we can construct a basis with Softplus functions, we will start with an intermediate basis whose basis elements $q_{n,k}$ are defined as follows using the parameterized Softplus function:
$q_{n,k}(x)=p_{a}(2^{n}x-k)-2p_{a}(2^{n}x-(k+\frac{1}{2}))+p_{a}(2^{n}x-(k+1))$ (19)
$a=a(n,k)=4\ln 2\cdot 2K\cdot 2^{2n+2}.$ (20)
We then have the following lemma:
Lemma 2.
The functions $\chi_{[0,1]}(x)$, $s_{1}(x)=x$ and $q_{n,k}(x)~{}(n\geq 0,0\leq k\leq 2^{n}-1)$ form a Schauder basis for $C[0,1]$.
Proof.
Let us apply the stability property in Theorem 5 with $X=C[0,1]$ to perturb the ReLU basis $t_{n,k}$. Let $K$ denote the basis constant for the ReLU basis and note that the ReLU basis is a normalized basis with $||t_{n,k}||=1$ as required by Theorem 5.
We will perturb the ReLU basis elements $t_{n,k}$ using a parameterized Softplus function $p_{a}(x)$ instead of the ReLU function $r(x)$ to obtain our new basis elements $q_{n,k}$. As we show below, increasing $a$ as a suitable multiple of $2^{2n+2}$, will ensure that the individual perturbations get smaller, and that the total perturbation across all basis elements remains small.
A simple analysis of the parameterized Softplus function $p_{a}(x)$ (6) shows that $|p_{a}(x)-r(x)|$ attains its maximum at $x=0$, and this maximum value is $p_{a}(0)=\frac{\ln 2}{a}$. This maximum value drops as $\frac{1}{a}$ when we increase the sharpness parameter $a$.
We further observe that for $a$ fixed, $\sup|p_{a}(2^{n}x-k)-r(2^{n}x-k)|=\sup|p_{a}(x)-r(x)|=\frac{\ln 2}{a}$. That is, the supremum of the difference between $p_{a}$ and $r$ does not change when the function parameters are scaled and shifted. Since $\sup_{x}|p_{a}(x)-r(x)|$ equals $||p_{a}-r||$, we can see that the norm of the perturbation $||p_{a}-r||$ remains unchanged for a given $a$ even when the function parameters for $p_{a}(x)$ and $r(x)$ are scaled and shifted. This gives us the following bound on the perturbation error between individual basis elements $t_{n,k}$ and $q_{n,k}$:
$||t_{n,k}-q_{n,k}||$ $=\sup|(r(2^{n}x-k)-2r(2^{n}x-(k+\frac{1}{2}))+r(2^{n}x-(k+1)))$
$\quad\quad-(p_{a}(2^{n}x-k)-2p_{a}(2^{n}x-(k+\frac{1}{2}))+p_{a}(2^{n}x-(k+1)))|$
$\leq\sup|r(2^{n}x-k)-p_{a}(2^{n}x-k)|+2|r(2^{n}x-(k+\frac{1}{2}))-p_{a}(2^{n}x-(k+\frac{1}{2}))|$
$\quad\quad+|r(2^{n}x-(k+1))-p_{a}(2^{n}x-(k+1))|$
$=\sup|r(x)-p_{a}(x)|+2|r(x)-p_{a}(x)|+|r(x)-p_{a}(x)|$
$=4\sup|r(x)-p_{a}(x)|$
$=\frac{4\ln 2}{a}.$
The total perturbation error $\Delta$ across all basis elements can now be bounded as:
$\Delta$ $=\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}||t_{n,k}-q_{n,k}||$
$\leq\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\frac{4\ln 2}{a}$
$\leq\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\frac{1}{2K}\frac{1}{2^{n+2}}\frac{1}{2^{n}}$
$=\frac{1}{2K}\sum_{n=0}^{\infty}\frac{1}{2^{n+2}}$
$<\frac{1}{2K}$
Since the total perturbation $\Delta<\frac{1}{2K}$, we conclude that $q_{n,k}$ is a Schauder basis for $C[0,1]$. This proves Lemma 2. ∎
Notice that $q_{n,k}$ form a smooth basis, but the elements of this basis are not scaled and shifted versions of a single mother function. The sharpness parameter destroys this scale-shift property of the basis. We also note that this construction cannot be based on sigmoidal functions like $\tanh$ since they lack a sharpness parameter to control the supnorm error, and to thus bound the total perturbation.
We will now prove Theorem 3 establishing that $p_{a}(x)$ forms a basis. We mimic the proof of Theorem 1 though we make use of the basis elements $q_{n,k}$ defined above instead of the original Schauder basis elements $s_{n,k}$.
Proof.
We will start with a basis expansion using $q_{n,k}$ denoting the coefficient functionals as $\beta_{j}$ and $\beta_{n,k}$:
$f$ $=\beta_{0}\chi_{[0,1]}+\beta_{1}s_{1}+\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\beta_{n,k}q_{n,k}$
$=\beta_{0}\chi_{[0,1]}+\beta_{1}s_{1}+\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\beta_{n,k}\{p_{a}(2^{n}x-k)-2p_{a}(2^{n}x-(k+\frac{1}{2}))+p_{a}(2^{n}x-(k+1))\}$
$=\beta_{0}\chi_{[0,1]}+\beta_{1}s_{1}+\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\{\beta_{n,k}p_{a}(2^{n}x-k)-2\beta_{n,k}p_{a}(2^{n}x-(k+\frac{1}{2}))+\beta_{n,k}p_{a}(2^{n}x-(k+1))\}$
$=\beta_{0}\chi_{[0,1]}+\beta_{1}s_{1}+\sum_{n=0}^{\infty}\{\beta_{n,0}p_{a}(2^{n}x)+\sum_{k=1}^{2^{n}-1}\{(\beta_{n,k}+\beta_{n,k-1})p_{a}(2^{n}x-k)-2\beta_{n,k}p_{a}(2^{n}x-(k+\frac{1}{2}))\}\}$ ($\ast$)
$=\beta_{0}\chi_{[0,1]}+\beta_{1}s_{1}+\sum_{n=0}^{\infty}\sum_{k=0}^{2^{n}-1}\{(\beta_{n,k}+\beta_{n,k-1})p_{a}(2^{n}x-k)-2\beta_{n,k}p_{a}(2^{n}x-(k+\frac{1}{2}))\}$
where, for convenience, we have set $\beta_{n,-1}=0$ in the last step. Like in the earlier proof, we note that ($\ast$4) preserves conditional convergence. Finally, given that the coefficient functionals $\beta_{j}$ and $\beta_{n,k}$ are bounded, the coefficients of the Softplus expansion, namely, $\gamma_{0}=\beta_{0}$, $\gamma_{1}=\beta_{1}$, $\gamma_{n,k}=(\beta_{n,k}+\beta_{n,k-1})$ and $\psi_{n,k}=-2\beta_{n,k}$ are all bounded as well. This establishes that the sequence of functions $\chi_{[0,1]}(x)$, $s_{1}(x)=x$, $p_{a}(2^{n}x-k)$ and $p_{a}(2^{n}x-(k+\frac{1}{2})$ form a Schauder basis, and concludes the proof of Theorem 3. ∎
As a simple corollary, we construct a sigmoidal basis based on Softplus functions. The construction of $u_{n,k}$ mimics the construction of the first discrete derivative $d_{n,k}$ in Theorem 2.
Theorem 6.
The functions $u_{n,k}(x)$ defined to be $p_{a(n,k)}(2^{n}x-k)-p_{a(n,k)}(2^{n}x-(k+\frac{1}{2}))$, with $a(n,k)=4\ln 2\cdot 2K\cdot 2^{2n+2}$, for $n\geq 0,0\leq k\leq 2^{n}-1$, along with the auxiliary functions $\chi_{[0,1]}(x)$ and $s_{1}(x)=x$ form a sigmoidal Schauder basis for $C[0,1]$
Observe that $u_{n,k}$ are indeed sigmoidal functions since $\lim_{x\to-\infty}u_{n,k}(x)$ equals 00, and $\lim_{x\to\infty}u_{n,k}(x)$ equals $1$. We skip the proof of the basis expansion as it follows the same lines as Theorem 2 except we use $q_{n,k}$ for the initial basis expansion instead of $t_{n,k}$.
5 Conclusion
We constructed four Schauder bases for $C[0,1]$, one using the ReLU function (Theorem 1), another using the Softplus function (Theorem 3), and two more using sigmoidal versions of the above (Theorem 2 and Theorem 6).
In terms of future work, we wonder if scaled and shifted bases for $C[0,1]$ are possible using smooth functions. In particular, we pose the following question: does there exist a smooth function $\sigma:[0,1]\to\mathbb{R}$ such that $\sigma_{n,k}(x)=\sigma(2^{n}x-k)$ forms a basis for $C[0,1]$printbibliography